OpenStudy (hpfan101):

f(x)=(5x^6 + 2x^3)^4

OpenStudy (hpfan101):

I did the chain rule and got: \[f'(x)=4(5x^6+2x^3)^3\times(30x^5+6x^2)\] But how would we simplify this?

OpenStudy (hpfan101):

My textbook is giving me an answer that I don't know how it got to that.

OpenStudy (anonymous):

What is the answer in your textbook?

OpenStudy (hpfan101):

It's: \[f'(x)=24x^11(5x^3+2)^3(5x^3+1)\]

OpenStudy (hpfan101):

24x^11

OpenStudy (anonymous):

ok We have \[f'(x)=4(5x^6+2x^3)^3(30x^5+6x^2)\] I want you to notice the first parenthesis \[(5x^6+2x^3)\] What can you factor out from this?

OpenStudy (hpfan101):

We can factor out an x^3

OpenStudy (anonymous):

exactly! \[(5x^6+2x^3)=x^3(5x^2+2)\] But!! remember, all this is actually CUBED so we have \[(5x^6+2x^3)^{3}=(x^3(5x^2+2))^3=??\]

OpenStudy (anonymous):

Distribute the cube

OpenStudy (anonymous):

don't expand it though, that will make it messy!

OpenStudy (hpfan101):

Ok so would it be: x^9(5x^3+2)^3? I'm not exactly sure since I usually expand whenever I cube something.

OpenStudy (anonymous):

exactly! so we have \[f'(x)=4x^9(5x^2+2)(30x^5+6x^2)\] Now factor our as much as u can from the 2nd parenthesis

OpenStudy (anonymous):

sorry theres a cube on the 5x^2+2

OpenStudy (hpfan101):

Oh okay! So the second paranthesis would be: 6x^2(5x^3+1)

OpenStudy (hpfan101):

And then we just multiply the 4x^9 and 6x^2 together to get 24x^11

OpenStudy (anonymous):

yep!\[f'(x)=4x^9(5x^2+2)^3.6x^2(5x^3+1)\] Multiply them together

OpenStudy (anonymous):

yup!!

OpenStudy (hpfan101):

Cool! And we get the same answer like in the book! Thank you so much! :)

OpenStudy (anonymous):

you're welcome!